# Estimation of Optimal Causal Bounds via Covariate-Assisted Optimal Transport

### v1

@@ Abstract
We study the estimation of causal estimand involving the joint distribution of treatment and control outcomes for a single unit. In typical causal inference settings, it is impossible to observe both outcomes simultaneously, which places our estimation within the domain of partial identification (PI). Pre-treatment covariates can substantially reduce estimation uncertainty by shrinking the partially identified set. Recently, it was shown that covariate-assisted PI sets can be characterized through conditional optimal transport (COT) problems. However, finite-sample estimation of COT poses significant challenges, primarily because, as we explain, the COT functional is discontinuous under the weak topology, rendering the direct plug-in estimator inconsistent. To address this issue, existing literature relies on relaxations or indirect methods involving the estimation of non-parametric nuisance statistics. In this work, we demonstrate the continuity of the COT functional under a stronger topology induced by the adapted Wasserstein distance. Leveraging this result, we propose a direct, consistent, non-parametric estimator for COT value that avoids nuisance parameter estimation. We derive the convergence rate for our estimator and validate its effectiveness through comprehensive simulations, demonstrating its improved performance compared to existing approaches.
Keywords: Conditional optimal transport, adapted Wasserstein distance, partial identification, causal inference.
@@ 1 Introduction
The potential outcome model [rubin1974estimating, imbens2015causal] is prevalent in causal inference, where each unit is associated with two potential outcomes: one under treatment and one under control. Since the two potential outcomes are never observed simultaneously, their joint distribution is not identified, making the causal estimands that depend on the joint distribution only partially identifiable. Optimal transport [villani2009optimal], which minimizes an expected cost over joint distributions respecting the marginals, has been used to recover partial identification (PI) sets of causal estimands [gao2024bridging]. When pre-treatment covariates are available, they can often be used to reduce uncertainty about the joint distribution of potential outcomes, and produce smaller, more informative PI sets. Conditional optimal transport (COT) has been used to characterize covariate-assisted PI sets [lin2025tightening, fan2025partial] by finding the optimal joint distribution preserving the outcome-covariate distributions.
@@
Despite the potential to yield more meaningful covariate-assisted PI sets, COT poses more substantial statistical challenges compared to OT, especially with continuous covariates. First, it is rare to observe any continuous covariate value in both treatment and control groups, leaving the conditional marginal distribution ill-defined at that value for at least one group. Second, even if each covariate value is associated with both a treated and a control unit (such as in twin studies), there is typically only one observation per covariate value in each group. The plug-in estimator of COT value, which uses point masses to estimate the conditional marginal distributions, may not converge to the true COT value as the sample size goes to infinity ([chemseddine2024conditional] and Example 3.1). This issue suggests that the COT functional may not be continuous under the weak topology, rendering the COT value estimation more challenging than standard OT problems.
@@
In this paper, we demonstrate the continuity of the COT functional under a stronger topology induced by the adapted Wasserstein distance, which requires uniform convergence of the outcome’s conditional distribution at each covariate value. Leveraging the continuity result, we propose a direct, consistent, non-parametric estimator for the COT value. Our estimator is constructed using an adapted empirical distribution, which discretizes the covariate space by assigning observed covariates to a finite number of cells. The convergence of the adapted empirical distribution under the adapted Wasserstein distance ensures the consistency of our estimator, and we further establish its finite-sample error bounds. Compared to existing estimators for COT value that rely on relaxations of the original COT formulation [manupriya2024consistent, lin2025tightening], or utilize indirect approaches that require estimating nuisance functions [ji2023model], or resort to heuristic procedures without statistical guarantees [tabak2021data, bunne2022supervised, hosseini2023conditional], our proposed COT value estimator is both direct and statistically consistent.
(a)           (b)           (c)
Figure 1: Illustration of our COT value estimator. For the observed covariates in the treatment and control group (Panel (a)), we first discretize the covariate space by assigning observed covariates to a finite set of cells (Panel (b)). We then match the discretized control and treatment covariates (Panel (c)), compute the OT value for each pair of discretized values, and aggregate.
@@
  1. (a)
    We establish the COT formulation the covariate-assisted PI sets of a class of causal estimands (Theorem 3.1). We prove the continuity of the COT functional under the topology induced by the adapted Wasserstein distance (Theorem 3.2).
  2. (b)
    Building on (a), under Bernoulli design, we propose a direct nonparametric estimator for the covariate-assisted PI set (or COT value), prove its consistency (Theorem 4.1), and further establish the finite-sample convergence rate (Theorem 4.2). We then extend the result to unconfounded designs with covariate-dependent treatment assignment (Theorem 4.3). In addition to its relevance for causal inference (through PI sets), our work advances the literature on COT by providing a consistent estimator for the COT value with provable convergence rate.
  3. (c)
    We validate the estimator’s effectiveness and robustness through comprehensive empirical studies (Section 5).
Organization. The rest of the paper is structured as follows. In Section 2, we introduce the problem setting along with preliminaries on COT. Section 3 provides basic properties of COT values, which serve as building blocks for our proposal. Section 4 presents the construction of our COT value estimator and the main theoretical results concerning its finite-sample performance. In Section 5, we provide simulation studies comparing our method with existing approaches. Technical proofs are deferred to the Appendix.
Notation. We denote $[n]=\{1,...,n\},n\wedge m=\min(n,m),n\vee m=\max(n,m)$. We denote by $\mathcal{P}(\Omega)$ the set of probability measures on $\Omega$. We use $\mu_{Y,Z}$ to denote the probability distribution (and its associated measure on $\mathcal{Y}\times\mathcal{Z}$) of $(Y,Z)$ under $\mu$, with $\mu_{Z}$ denoting the marginal distribution of $Z$, and $\mu^{z}_{Y}$ the conditional distribution of $Y$ given $Z=z$. By $\mu=\mu_{Z}\otimes\mu_{Y}^{Z}$, we mean that for any measurable function $g$, $\int g{\rm d}\mu=\int_{\mathcal{Z}}\int_{\mathcal{Y}}g(z,y){\rm d}\mu_{Y}^{z}(y){\rm d}\mu_{Z}(z)$. We denote the set of the joint couplings $\pi$ with marginals $\mu,\nu\in\mathcal{P}(\mathcal{X})$ ($\mathcal{X}$ may be $\mathcal{Y},\mathcal{Z}$ or $\mathcal{Y}\times\mathcal{Z}$) by
@@
Below, we use covariate-assisted PI sets and PI sets exchangeably.
@@ 2.1 Potential Outcome Model
We consider the potential outcome model with a binary treatment [rubin1974estimating]. Suppose there are $n$ units. Each unit is associated with two potential outcomes, $Y_{i}(0)$ and $Y_{i}(1)\in\mathcal{Y}\subseteq\mathbb{R}^{d_{Y}}$, $i\in[n]$, $d_{Y}\geq 1$, $\mathcal{Y}$ a convex closed set. Here $Y_{i}(0)$ is the outcome under control and $Y_{i}(1)$ is the outcome under treatment [^1] We allow the potential outcomes to be multivariate to accommodate trials with multiple endpoints, experiments with a primary and secondary outcome, or policy evaluations involving effectiveness and cost.. Each unit receives a binary treatment assignment $W_{i}\in\{0,1\}$, where $W_{i}=1$ indicates treatment and $W_{i}=0$ indicates control. For unit $i$, only the potential outcome corresponding to the received treatment level, i.e., $Y_{i}(W_{i})$, is observed.
@@
We impose the following standard assumptions [imbens2015causal] for the potential outcome model.
Assumption 2.1 (Stable Unit Treatment Value Assumption (SUTVA)).
@@
To enable the asymptotic analysis, we adopt the super-population model [imbens2015causal] justified by Assp. 2.1, 2.2,
@@
where $\mu$ denotes some unknown joint distribution. The following result shows that, under the super-population model with unconfounded treatment assignment, the outcome-covariate distributions $(Y_{i}(1),Z_{i})$ and $(Y_{i}(0),Z_{i})$ are identifiable from the observed data. The outcome-covariate distributions will be later used in the COT formulation of the PI sets.
Proposition 2.1 (Identifiable marginals).
Under Assp. 2.1 - 2.3, then for $k=0,1$,
$\mathbb{P}\left(Y_{i}(k)=y,Z_{i}=z\right)$ $=w_{k}(z)\ \mathbb{P}\left(Y_{i}=y,Z_{i}=z\mid W_{i}=k\right),$ (2)
where $w_{1}(z)={\bar{e}}/{e(z)}$, $w_{0}(z)={(1-\bar{e})}/{(1-e(z))}$, $\bar{e}:=\int e(z){\rm d}\mu_{z}(z)$.
@@ 2.2 PI Sets of Causal Estimands
Here $h:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$ is some pre-specified objective (cost) function.
A fundamental challenge in causal inference is that $Y_{i}(0)$ and $Y_{i}(1)$ cannot be observed simultaneously for the same unit, making the joint distribution $\mu_{Y(0),Y(1)}$ inherently unidentifiable. As a result, for causal estimands depending on the joint distribution, their values are not identifiable from observed data [^2] Non-identifiability refers to the fact that two different joint distributions can yield different values of the causal estimand while producing the same distribution over observable quantities. . This issue is commonly known as partial identification; see Examples 2.1, 2.2 for illustration.
Example 2.1 (Variance of treatment effects).
For $h(y_{0},y_{1})=(y_{0}-y_{1})^{2}$, the corresponding estimand represents the variance of the treatment effect [^3] This estimand can be used to construct a confidence interval for the average treatment effect $\mathbb{E}[Y(1)-Y(0)]$ in the classic Neymanian framework [neyman1923application]. and is only partially identifiable [lin2025tightening].
Example 2.2 (Distributional treatment effects).
For $h(y_{0},y_{1})=\mathbb{I}(y_{1}-y_{0}\leq t)$, $t\in\mathbb{R}$, the resulting estimand is referred to as the distributional treatment effect and is only partially identifiable [williamson1990probabilistic, fan2010sharp, fava2024predicting].
@@
$\Pi_{\textup{c}}\triangleq\left\{\pi\in\mathcal{P}(\mathcal{Y}^{2}\times\mathcal{Z}):\pi_{Y(0),Z}=\mu_{Y(0),Z},\ \pi_{Y(1),Z}=\mu_{Y(1),Z}\right\}.$ (3)
@@
Here we use the fact that the coupling set $\Pi_{\textup{c}}$ defined in (3) is convex, the resulting PI set (4) is a convex subset of $\mathbb{R}$, and thus an interval [^4] The interval may be unbounded and contain $\infty$- ∞ or +$\infty$+ ∞.. To construct this PI interval, it suffices to compute its lower and upper bounds, starting with
@@
We conclude this section with a discussion of related work on PI sets. Partial identification has been extensively studied in causal inference and econometrics (see [kline2023recent] for a review), and we review three particularly relevant threads. First, copula models [nelsen2006introduction], and particularly the Fréchet–Hoeffding bounds [heckman1997making, manski1997mixing, fan2010sharp, kawakami2025moments], have been used to construct PI sets. However, these bounds are limited to univariate responses, only pointwise sharp in settings with more than two margins, and do not necessarily translate into bounds on causal estimands based on an objective function $h$. The second thread of related works explicitly employs the optimal transport to address the partial identification problem [galichon2018optimal], but not under the potential outcome model. Third, [ji2023model] address the partial identification problem using the COT framework, focusing on its dual formulation. Their method solves for the optimal dual function at each observed covariate value, and is thus indirect (depending on estimated nuisance functions) and sensitive to nuisance estimation error. In this paper, we estimate the covariate-assisted PI set by directly investigating the primal formulation of a COT problem.
@@ 2.3 Conditional Optimal Transport
We begin by reviewing optimal transport (OT), and then extend the discussion to conditional optimal transport (COT).
OT has a rich literature [villani2009optimal] with broad applications in machine learning and statistics. In particular, OT has recently emerged as a powerful tool in causal inference [black2020fliptest, charpentier2023optimal, de2024transport, torous2024optimal]. In an OT problem, the goal is to couple two distributions in a way that minimizes a specified cost function (e.g., $h$) while preserving their marginal distributions. Formally, the OT distance is defined as
@@
Compared to OT, COT aligns conditional measures across values of the covariates, which has seen various applications, such as conditional sampling [wang2023efficient, hosseini2023conditional, baptista2024representation], domain adaptation [rakotomamonjy2022optimal], and Bayesian flow matching [kerrigan2024dynamic, chemseddine2024conditional]. In this paper, we focus on the natural connection between COT and the PI problem when covariates are available. Particularly, recall that in the PI set problem with $Z_{i}$, it becomes essential to preserve the outcome-covariate distributions of $(Y_{i}(0),Z_{i})$ and $(Y_{i}(1),Z_{i})$, where $Z_{i}$ is shared across both marginals. The corresponding covariate-assisted lower bound is defined in Eq. (5), which, as shown below, is the optimal objective value of a COT problem.
Proposition 2.2 (Recursive formulation).
@@
We conclude this section with the topology for the COT functional, which motivates our finite-sample estimator. As we will demonstrate in Section 3, a key distinction between OT and COT as functionals lies in their topological properties: while the OT functional is continuous with respect to the outcome marginals under the weak topology (metrized by $\mathds{W}_{1}$), the COT functional $V_{\textup{c}}$ is discontinuous with respect to the outcome-covariate marginals under the weak topology. However, as shown in Theorem 3.2, $V_{\textup{c}}$ becomes continuous on the set of $\mu$ with Lipschitz kernel (see Assp. 3.4 below) under a stronger topology induced by the so-called adapted Wasserstein distance.
Definition 2.1 (Adapted Wasserstein distance).

### v2

# Causal Partial Identification via Conditional Optimal Transport

@@ Abstract
We study the estimation of causal estimand involving the joint distribution of treatment and control outcomes for a single unit. In typical causal inference settings, it is impossible to observe both outcomes simultaneously, which places our estimation within the domain of partial identification (PI). Pre-treatment covariates can substantially reduce estimation uncertainty by shrinking the partially identified set. Recent work has shown that covariate-assisted PI sets can be characterized through conditional optimal transport (COT) problems. However, finite-sample estimation of COT poses significant challenges, primarily because the COT functional is discontinuous under the weak topology, rendering the direct plug-in estimator inconsistent. To address this issue, existing literature relies on relaxations or indirect methods involving the estimation of non-parametric nuisance statistics. In this work, we demonstrate the continuity of the COT functional under a stronger topology induced by the adapted Wasserstein distance. Leveraging this result, we propose a direct, consistent, non-parametric estimator for COT value that avoids nuisance parameter estimation. We derive the convergence rate for our estimator and validate its effectiveness through comprehensive simulations, demonstrating its improved performance compared to existing approaches.
@@ 1 Introduction
The potential outcome model [44, 27] is prevalent in causal inference, where each unit is associated with two potential outcomes: one under treatment and one under control. Since the two potential outcomes are never observed simultaneously, their joint distribution is not identified, making the causal estimands that depend on the joint distribution only partially identifiable. Optimal transport [49], which minimizes an expected cost over joint distributions respecting the marginals, has been used to recover partial identification (PI) sets of causal estimands [24]. When pre-treatment covariates are available, they can often be used to reduce uncertainty about the joint distribution of potential outcomes, and produce smaller, more informative PI sets. Conditional optimal transport (COT) has been used to characterize covariate-assisted PI sets [28, 21, 35, 6] by finding the optimal joint distribution preserving the outcome-covariate distributions.
@@
Despite the potential to yield more meaningful covariate-assisted PI sets, COT poses more substantial statistical challenges compared to OT, especially with continuous covariates. First, it is rare to observe a shared continuous covariate value in both treatment and control groups, leaving the conditional marginal distribution ill-defined at that value for at least one group. Specifically, when conditioning on an exact covariate value, there will typically be only a single unit in the sample that attains that value, with matched or repeated observations being unlikely. Second, even if each covariate value is associated with both a treated and a control unit (such as in twin studies), there is typically only one observation per covariate value in each group. The plug-in estimator of COT value, which uses point masses to estimate the conditional marginal distributions, may not converge to the true COT value as the sample size goes to infinity ([16] and Example 2). This issue suggests that the COT functional may not be continuous under the weak topology, rendering the COT value estimation more challenging than standard OT problems.
@@
In this paper, we demonstrate the continuity of the COT functional under a stronger topology induced by the adapted Wasserstein distance, which requires uniform convergence of the outcome’s conditional distribution at each covariate value. Leveraging the continuity result, we propose a direct, consistent, non-parametric estimator for the COT value. Our estimator is constructed using an adapted empirical distribution, which discretizes the covariate space by assigning observed covariates to a finite number of cells. The convergence of the adapted empirical distribution under the adapted Wasserstein distance ensures the consistency of our estimator, and we further establish its finite-sample error bounds. Compared to existing estimators for COT value that rely on relaxations of the original COT formulation [38, 35], or utilize indirect approaches that require estimating nuisance functions [28], or resort to heuristic procedures without statistical guarantees [46, 12, 26], our proposed COT value estimator is direct and statistically consistent.
@@
  1. 1.
    We establish the COT formulation for the covariate-assisted PI sets of a class of causal estimands (Theorem 1). We prove the continuity of the COT functional under the topology induced by the adapted Wasserstein distance (Theorem 2).
  2. 2.
    Building on (a), under Bernoulli design, we propose a direct nonparametric estimator for the covariate-assisted PI set, prove its consistency (Theorem 3), and further establish the finite-sample convergence rate (Theorem 4) together with a robustness guarantee under covariate’s distributional shift (Theorem 5). We then extend the result to unconfounded designs with covariate-dependent treatment assignment (Theorem 6). In addition to its relevance for causal inference, our work advances the literature on COT by providing a consistent non-parametric estimator for the COT value with provable convergence rate.
  3. 3.
    We validate the estimator’s effectiveness through comprehensive empirical studies (Section 5).
Organization. The rest of the paper is structured as follows. Section 2 introduces the problem setting. Section 3 provides basic properties of COT values, which serve as building blocks for our proposal. Section 4 presents the construction of our COT value estimator and the main theoretical results on its finite-sample performance. In Section 5, we provide simulation studies comparing our method with existing approaches. Technical proofs are deferred to the appendix.
Notation. We use $Z(w)$ to denote the variable $Z$ observed within the control ($w=0$) and treatment ($w=1$) groups. We denote $n]=\{1,...,n\},n\wedge m=\min(n,m),n\vee m=\max(n,m)$. We denote by $\mathcal{P}(\Omega)$ the set of probability measures on $\Omega$. We use $\mu_{Y,Z}$ to denote the probability distribution (and its associated measure on $\mathcal{Y}\times\mathcal{Z}$) of $(Y,Z)$ under $\mu$, with $\mu_{Z}$ denoting the marginal distribution of $Z$, and $\mu^{z}_{Y}$ the conditional distribution of $Y$ given $Z=z$. By $\mu=\mu_{Z}\otimes\mu_{Y}^{Z}$, we mean that for any measurable function $g$, $\int g{\rm d}\mu=\int_{\mathcal{Z}}\int_{\mathcal{Y}}g(z,y){\rm d}\mu_{Y}^{z}(y){\rm d}\mu_{Z}(z)$. We denote the set of the joint couplings $\pi$ with marginals $\mu,\nu\in\mathcal{P}(\mathcal{X})$ ($\mathcal{X}$ may be $\mathcal{Y},\mathcal{Z}$ or $\mathcal{Y}\times\mathcal{Z}$) by
@@ 2.1 Potential Outcome Model
We consider the potential outcome model with a binary treatment [44]. Suppose there are $N$ units. Each unit is associated with two potential outcomes, $Y_{i}(0)$ and $Y_{i}(1)\in\mathcal{Y}\subseteq\mathbb{R}^{d_{Y}}$, $i\in[N]$, $d_{Y}\geq 1$, $\mathcal{Y}$ a convex closed set. Here $Y_{i}(0)$ is the outcome under control and $Y_{i}(1)$ is the outcome under treatment [^1] We allow the potential outcomes to be multivariate to accommodate trials with multiple endpoints, experiments with a primary and secondary outcome, or policy evaluations involving effectiveness and cost.. Each unit receives a binary treatment $W_{i}\in\{0,1\}$, where $W_{i}=1$ indicates treatment and $W_{i}=0$ indicates control. For unit $i$, only the outcome $Y_{i}(W_{i})$ is observed.
@@
We impose the following standard assumptions [27] for the potential outcome model.
Assumption 1 (Stable Unit Treatment Value Assumption (SUTVA)).
@@
To enable the asymptotic analysis, we adopt the super-population model [27] justified by Assp. 1, 2,
@@
where $\mu$ is the unknown joint distribution. The following result shows that, under the super-population model with unconfounded treatment assignment, the outcome-covariate distributions $(Y_{i}(1),Z_{i})$ and $(Y_{i}(0),Z_{i})$ are identifiable from the observed data. The outcome-covariate distributions will be later used in the COT formulation of the PI sets.
Proposition 1 (Identifiable marginals).
Under Assp. 1 - 3, for $w=0,1$,
$\mathbb{P}\left(Y_{i}(w)=y,Z_{i}=z\right)$
$=$~ $\tilde{e}_{w}(z)\ \mathbb{P}\left(Y_{i}(W_{i})=y,Z_{i}=z\mid W_{i}=w\right),$ (2)
where $\tilde{e}_{w}(z)=\frac{(1-w)-\bar{e}}{(1-w)-e(z)}$, $\bar{e}\triangleq\int e(z){\rm d}\mu_{z}(z)$.
@@ 2.2 PI Sets of Causal Estimands
Here, $h:\mathcal{Y}^{2}\rightarrow\mathbb{R}$ is a pre-specified objective function.
A fundamental challenge in causal inference is that $Y_{i}(0)$ and $Y_{i}(1)$ cannot be observed simultaneously for the same unit, making the joint distribution $\mu_{Y(0),Y(1)}$ inherently unidentifiable. As a result, for causal estimands depending on the joint distribution, their values are not identifiable from observed data [^2] Non-identifiability refers to the fact that two different joint distributions can yield different values of the causal estimand while producing the same distribution over observable quantities. . This issue is commonly known as partial identification.
Example 1 (Variance of treatment effects).
For $h(y_{0},y_{1})=(y_{0}-y_{1})^{2}$, the corresponding estimand represents the variance of the treatment effect, which can be used to construct a confidence interval for the average treatment effect $\mathbb{E}[Y(1)-Y(0)]$ in the classic Neymanian framework [45]. This estimand is only partially identifiable and has been studied in [2, 7].
@@
$\Pi_{\textup{c}}\triangleq\left\{\pi\in\mathcal{P}(\mathcal{Y}^{2}\times\mathcal{Z}):\pi_{Y(w),Z}=\mu_{Y(w),Z}\ \ w=0,1\right\}.$
@@
Here, we use the fact that the set $\Pi_{\textup{c}}$ is convex, the resulting PI set (3) is a convex subset of $\mathbb{R}$, and thus an interval. To construct this interval, it suffices to compute its lower and upper bounds, starting with
@@
Note that, for certain objective functions $h$ (including Example 1), Hoeffding–type results yield a closed-form optimizer for (4) (see, e.g., [47]). In contrast, we develop estimators for (4) that apply to broad classes of smooth functions $h$.
@@ 2.3 Conditional Optimal Transport
We first review optimal transport (OT), followed by a discussion of conditional optimal transport (COT).
OT has a rich literature [49] with broad applications in machine learning and statistics. In particular, OT has recently emerged as a powerful tool in causal inference [10, 15, 19, 48]. In an OT problem, the goal is to couple two distributions in a way that minimizes a specified cost function (e.g., $h$) while preserving their marginal distributions. The OT distance is defined as
@@
Compared to OT, COT aligns conditional measures across values of the covariates, which has seen various applications, such as conditional sampling [50, 26, 8], domain adaptation [42], and Bayesian flow matching [32, 16]. In this paper, we focus on the natural connection between COT and the PI problem when covariates are available. Particularly, recall that in the PI set problem with $Z_{i}$, it becomes essential to preserve the outcome-covariate distributions of $(Y_{i}(0),Z_{i})$ and $(Y_{i}(1),Z_{i})$, where $Z_{i}$ is shared across both marginals. The corresponding covariate-assisted lower bound is defined in Eq. (4), which, as shown below, is the optimal objective value of a COT problem.
Proposition 2 (Recursive formulation).
@@
2.4 Adapted Wasserstein Distance and Its Topology
We conclude this section by introducing the topology that is natural for the COT functional and that motivates our finite-sample estimator in Section 4.
As we will demonstrate in Section 3, a key distinction between OT and COT as functionals lies in their topological properties. While the classical OT functional is continuous with respect to the outcome marginals under the weak topology (metrized by $\mathds{W}_{1}$), the COT functional $V_{\textup{c}}$ is generally discontinuous with respect to the joint outcome–covariate marginals under the weak topology. This discrepancy arises because COT is inherently a conditional transport problem: its geometry is driven by transition kernels rather than full joint distributions.
A natural way to formalize this structure is to restrict the set of admissible couplings to those that respect the conditional structure of the marginals, i.e., to (bi-)causal or adapted couplings; see, e.g., [40, 34, 5]. From this perspective, the adapted Wasserstein distance can be understood as the optimal transport distance obtained when couplings are constrained to match conditional distributions. This interpretation is particularly natural in causal inference settings, where one aims to compare conditional outcome laws given covariates rather than arbitrary joint couplings of $(Y,Z)$ and $(Y^{\prime},Z^{\prime})$.
We therefore equip $\mathcal{P}(\mathcal{Y}\times\mathcal{Z})$ with the topology induced by the adapted Wasserstein distance.
Definition 1 (Adapted Wasserstein distance).